OpenStudy (anonymous):

solve log-5x=log(10-3x)

OpenStudy (anonymous):

when u find a log in the both sides but should have the same base and the same function ,u can cancel it and solve a normal equation -5x=10-3x -5x+3x=10 -2x=10 x=-5